On January 1, 2020, ABC Company leased equipment from ZZ Leasing, Inc. The terms of the lease require annual payments of $7,500 for ten years with the first payment due December 31, 2020. The interest rate on the lease is 3%. Assume that the lease qualifies as a capital lease. ABC Company uses the double-declining balance method to depreciate its assets. Calculate the book value of the leased asset at December 31, 2021.
Answer #1

Solution:

Fair value of leased asset = Present value of lease payments

= $7,500 * Cumulative PV factor at 3% for 10 periods

= $7,500 * 8.53020 = $63,977

Depreciation rate - SLM = 1/10 = 10%

Depreciation rate - DDB = 10%*2 = 20%

Depreciation for 2020 = $63,977*20% = $12,795

Book value at the beginning of 2021 = $63,977 - $12,795 = $51,182

Depreciation for 2021 = $51,182 * 20% = $10,236

Book value of leased asset at December 31, 2021 = $51,182 - $10,236

= $40,946